OpenStudy (sonyalee77):

Please Help !!! Medals!!! Which of the following are point-slope equations of the line going through (-4, 6) and (1, 2)?

OpenStudy (sonyalee77):

@Nateyboo

OpenStudy (anonymous):

What are the options?

OpenStudy (sonyalee77):

A. y + 4 = -4/5(x - 6) B. y - 6 = -4/5(x + 4) C. y - 6 = -5/4(x + 4) D. y - 2 = -4/5(x - 1) E. y + 2 = 4/5 (x - 1) F. y - 2 = -5/4(x - 1)

OpenStudy (sonyalee77):

@wolfe8

OpenStudy (sonyalee77):

@eliassaab

OpenStudy (anonymous):

(-4, 6) and (1, 2) \[ slope=\frac{6-2}{-4-1}=-\frac {4}{5} \]

OpenStudy (anonymous):

y-6=-4/5(x+4)

OpenStudy (anonymous):

or y-2=-4/5( x-1)

OpenStudy (anonymous):

Both are fine

OpenStudy (anonymous):

It seems B and D

OpenStudy (wolfe8):

I would plug in the values and see.

OpenStudy (anonymous):

@Sonyalee77 : I'll agree with eliassaab's answer.

OpenStudy (sonyalee77):

Thanx all i needed to know basically was the slope :P

OpenStudy (anonymous):

The slope will eliminate some possibilities and the the slope point formula will get you the two possible answers